40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 1 08:04:13 2018, Jon Richardson, Configuration, Electronics, PSL-Aux. Laser Phase-Locked Loop IMG_2553.jpgIMG_2551.jpgIMG_2552.jpgsrc_gouy_phase_v3.pdf
    Reply  Thu May 3 20:29:39 2018, gautam, Configuration, Electronics, PSL-Aux. Laser Phase-Locked Loop 
Message ID: 13813     Entry time: Thu May 3 20:29:39 2018     In reply to: 13802
Author: gautam 
Type: Configuration 
Category: Electronics 
Subject: PSL-Aux. Laser Phase-Locked Loop 

Some notes about the setup and work at the PSL table today, Jon can add to / correct me.

  • All equipment for the phase locking now sit on a cart that is on the west side of the MC beam tube, near ITMX chamber.
  • Cables have been routed through the space between the PSL enclosure and the optical table.
  • HEPA was turned up for this work, now it has been turned down to the nominal level of 30%.
  • Alignment into the PMC had degraded a bit - I tweaked it and now MC transmission is up at ~15600 which is a number I am used to. We still don't have a PMC transmission monitor since the slow ADC failure.
ELOG V3.1.3-